A bottle is heated with its mouth open from 15`C to 100`C. What fraction of air originally contained in the vessel is expelled?

Grade:11

2 Answers

Joana
7 years ago
Show tht
Laltesh Singh
6 years ago
As we know PV = nRT​From the above relation we know that temperature of gas is inversely proportional to the number of moles.n2/n1= 288/3731 - n2/n1 = 1 - 288/373n1-n2/n1 = 85/373n1-n2/n1 = 0.2278